OpenStudy (anonymous):

which mathematical property is demonstrated 5+3w+2=5+2+3w

OpenStudy (serenity74):

What do you think is the answer?

OpenStudy (serenity74):

And why do you think that's the answer?

OpenStudy (anonymous):

i don't know thats why i posted it...

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

Here are the properties of addition tell me which matches your problem :) 

OpenStudy (serenity74):

What are the diffferent types of properties? I can't just give you the answer mate.

OpenStudy (anonymous):

@pooja195 what do u mean

OpenStudy (anonymous):

and @Serenity74 u mean this comutative property of addition those ?

pooja195 (pooja195):

Looks at the different properties i showed you which relates to your question? Commutative property says that any addtion problem can be rearranged and you would still get the same answer ^would that be your answer? Yes or no? (im just going down the list)

OpenStudy (serenity74):

Yea that's what I would swing for.

OpenStudy (anonymous):

@pooja195 yes cuz its adding??

pooja195 (pooja195):

Yes \[\huge~\rm~\bf~5+3w+2=5+2+3w=\color{purple}{3w+7}\]

OpenStudy (anonymous):

yay

pooja195 (pooja195):

;-; it got cut it says 3w+7 Basically trying to say no matter what order you put those numbers/variables in they will equal the same

OpenStudy (anonymous):

ok but why is it not associative property of addition

pooja195 (pooja195):

That's used for grouping so